OpenStudy (anonymous):

tan^2x-1<0 solve for x

OpenStudy (anonymous):

\[\text{Reduce}\left[\text{Tan}[x]^2 - 1 < 0\right] \to -1<\text{Tan}[x]<1 \]

OpenStudy (anonymous):

You can always come back to twiddla...I can do the steps

OpenStudy (anonymous):

Let me know

OpenStudy (anonymous):

Actually, rotobey, your methodology is a bit off....You didn't really show the steps properly

OpenStudy (anonymous):

im confused still, can someone show me the steps

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

\[\text{Reduce}\left[(\text{Tan}[x])^2-1<0\right]= (-1<\text{Tan}[x]<1) \]is from Mathematica. Is the argument of Reduce correct?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

girl222, did you have an other problem you wanted solved?

OpenStudy (anonymous):

A complete solution from WolframAlpha.com is attached. It doesn't show the steps. Perhaps @heromiles or someone else will provide solution steps.